# FALCON: Fairness Learning via Contrastive Attention Approach to Continual Semantic Scene Understanding in Open World

### v1

@@ 3.3 Open-world Unknown Class Modeling
An ideal CSS approach must be able to model the unknown classes without supervision, especially in open-world environments [27, 52]. Prior studies have adopted the pseudo-label strategies [19, 4] based on the model predictions to assign labels for seen classes while unseen classes have been ignored, thus resulting in non-discriminative features. [27, 52] improved the background modeling by using an additional prototypical representation for unknown classes. However, these approaches consider different unknown classes as one (i.e., $N_{U}=1$) resulting in non-distinguished representations of different unknown classes. Thus, modeling function $\phi$ in Eqn. (2) without supervision of different unknown classes (i.e., $N_{U}>1$) is challenging.
Figure 4: The Enforcement Loss of Vanilla Contrastive Clustering $\mathcal{L}_{Cont}$ and Fairness Contrastive Clustering $\mathcal{L}^{\alpha}_{Cont}$ on Pascal VOC. Since $\mathcal{L}_{Cont}$ suffers severe biased between major and minor classes, its clusters of minor classes remain scattered. Meanwhile, our $\mathcal{L}^{\alpha}_{Cont}$ produces a more uniform loss among classes, thus promoting the fairness and compactness of clusters.
@@
Although modeling $\phi$ to determine the single feature $\mathbf{f}$ belonging to the cluster $\mathbf{c}$ is challenging, prior studies in clustering [35, 58, 57] have suggested that determine a set of features $\{\mathbf{f}^{t}_{i}\}_{i=1}^{M}$ belonging to cluster $\mathbf{c}$ should be easier. This derives from the fact that even though the feature representations of different classes are different, the distributions of features around its cluster (termed as Visual Grammar) in the feature space should be similar among classes or clusters. As a result, by learning the distribution of features and their clusters, the model $\phi$ can determine whether a feature belongs to a cluster. Then, by learning the model $\phi$ on prior known clusters and features, the knowledge of $\phi$ can be adaptively applied to unknown clusters. Fig. 5 illustrates our visual grammar model of the cluster distributions.
@@
Limitations of Prior Clustering Methods The traditional methods in clustering, e.g., KNN or density-based clustering [21], remain limited to noisy features leading to producing the incorrect cluster assignment. Meanwhile, the modern clustering methods, e.g., Graph Neural Networks (GNNs) [58, 57], require a large memory to build the affinity graph for clusters. In addition, GNNs often learn the local structures of graphs (or clusters) and accumulate them via the aggregation layers. Hence, the global structures of the clusters, i.e., visual grammar, are not well modeled by GNNs [35]. Therefore, to address these limitations, we introduced a new Attention-based Visual Grammar approach to efficiently model the distribution of features and their clusters via the self-attention mechanism [53].
@@
Remark 1: Given a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}$ where $\mathbf{f}^{\mathbf{c}}_{i}$ denotes the feature $\mathbf{f}_{i}$ belonging to the cluster $\mathbf{c}$, and $\forall i\in[1..M-1]:\cos(\mathbf{f}_{i}^{\mathbf{c}},\mathbf{c})\geq\cos(\mathbf{f}_{i+1}^{\mathbf{c}},\mathbf{c})$ the Visual Grammar of the cluster $\mathbf{c}$ parameterized by $\Theta$ can be defined as Eqn. (5).
@@
$\footnotesize\begin{split}&\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{f}^{\mathbf{c}}_{1},\mathbf{f}^{\mathbf{c}}_{2},...,\mathbf{f}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\\ =&\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{\Delta}^{\mathbf{c}}_{1},\mathbf{\Delta}^{\mathbf{c}}_{2},...,\mathbf{\Delta}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\end{split}$ (5)
@@
where $\Delta^{\mathbf{c}}_{i}=\mathbf{f}^{\mathbf{c}}_{i}-\mathbf{c}$. Eqn. (5) defines the visual grammar of the cluster by modeling the feature distribution of $\mathbf{f}_{i}^{\mathbf{c}}$ and its cluster center $\mathbf{c}$. Let $\phi:\mathbb{R}^{(M+1)\times D}\to[0,1]^{M}$ be a function receiving a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}_{i}\}_{i=1}^{M}$ $\left(\cos(\mathbf{f}_{i},\mathbf{c})\geq\cos(\mathbf{f}_{i+1},\mathbf{c})\right)$ to determine whether $\mathbf{f}_{i}$ belonging to $\mathbf{c}$, i.e., $\mathbf{u}=\phi(\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c})$ where $\mathbf{\Delta}_{i}=\mathbf{f}_{i}-\mathbf{c}$, $\mathbf{u}=[u_{1},u_{2},...,u_{M}]$ and $u_{i}=1$ denotes $\mathbf{f}_{i}$ belong to cluster $\mathbf{c}$ and vice versa. Hence, the visual grammar model in Eqn. (5) can be modeled by the network $\phi$ as follows with parameter $\Theta$ as follows:
@@
Eqn. (6) aims to model the distribution of features around its cluster by learning the correlation of relatively topological structures $\mathbf{\Delta}_{i}$ of features $\mathbf{f}_{i}$ around cluster $\mathbf{c}$. Then, based on knowledge of the cluster distribution, the model $\phi$ is able to determine whether a feature $\mathbf{f}_{i}$ belongs to cluster $\mathbf{c}$. Hence, it is essential that the model $\phi$ has the ability to exploit the correlation between features $\mathbf{f}_{i}$ and cluster $\mathbf{c}$ to learn the topological structure of visual grammar. Therefore, we adopt the self-attention mechanism [53, 35] to efficiently model these feature correlations. Particularly, the model $\phi$ is formed by $L_{\phi}$ blocks of self-attention as follows:
$\footnotesize\begin{split}\mathbf{z}_{0}&=\operatorname{LN([\mathbf{\Delta}_{1},...,\mathbf{\Delta}_{M},\mathbf{c}])}+\boldsymbol{\beta},\quad\mathbf{a}_{l}=\mathbf{z}_{l}+\operatorname{MHSA}(\mathbf{z}))\\ \mathbf{z}_{l+1}&=\mathbf{a}_{l}+\operatorname{MLP}(\operatorname{LN}(\mathbf{a}_{l})),\quad\quad\quad\quad\mathbf{u}=\operatorname{Proj}(\mathbf{z}_{L_{\phi}})\end{split}$ (7)
@@
Figure 5: The Proposed Visual Grammar Model.
Cluster Assignment via Visual Grammar Instead of assigning the clusters based on the model prediction [19, 4, 39] or nearest cluster [52, 27] that are less effective, the cluster assignment in our approach will be performed by the visual grammar model, i.e., the visual grammar model will consider the $M$ closest features around cluster $\mathbf{c}$ to assign the cluster for these features. Then, the cluster assignments are used to compute our Fairness Contrastive Clustering loss. In addition, following common practices [19, 4, 39], we improve background shift modeling by using the cluster assignments of features as the pseudo labels of pixels.
Unknown Cluster Initialization Prior work [52, 27] initialized a single unknown cluster ($N_{U}=1$), thus resulting in producing non-discriminative class-wise features. However, there should be more than a single unknown cluster ($N_{U}>1$) to produce discriminative features for different unknown classes. Therefore, our approach first initializes a list of potential unknown clusters at each learning step via DB-SCAN [21] on the features of unknown classes extracted by the current CSS model. For the new known class $\mathcal{C}^{t}$, we initialize these clusters based on the mean of their feature representations. Meanwhile, the clusters of known classes learned in previous steps are maintained.
@@ 3.4 Continual Learning Procedure
Fig. 3 illustrates the training procedure of our continual learning approach. At each learning step $t$, the CSS model $F$ with $\theta_{t}$ is trained with the Fairness Contrastive Clustering loss defined in Eqn. (4) and the previous visual grammar model $\phi$ with $\Theta_{t-1}$. In addition, we introduce a cluster regularizer $\mathcal{R}_{C}$ to avoid the clusters of different classes collapsing into a single cluster. Therefore, the entire CSS learning objective in our approach can be formed as:
$\scriptsize\begin{split}\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}}\Big{[}\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\sum_{\mathbf{c}_{i}}\mathcal{L}^{\alpha}_{Cont}\left(\mathbf{F}^{t},\mathbf{c}_{i}\right)+\lambda_{C}\mathcal{R}_{C}(\mathbf{c})\Big{]}\end{split}$ (8)
@@
Training Procedure of Visual Grammar Model At CSS learning step $t$, we adopt the visual grammar model trained on the previous learning step, i.e., $\phi$ with $\Theta_{t-1}$, to perform the cluster assignment for the contrastive clustering loss defined in Eqn. (2). Then, the visual grammar model at learning step $t$, i.e., $\phi$ with $\Theta_{t}$, will be learned (initialized from $\Theta_{t-1}$) on the features extracted from the dataset and the set of known clusters $\mathbf{c}$ up to the current learning step. Following [35], we sample a center $\mathbf{c}$ from the known clusters and its $M$ closest features to train the visual grammar model.
Initial Visual Grammar Model At the first learning step $t=1$, since no clusters have been learned at initial, the visual grammar model $\phi$ with $\Theta_{0}$ is not available. However, as common practices in CSS [19, 4, 5], the segmentation model is typically trained from a pre-trained backbone on ImageNet [17]. As a result, the features extracted at the first learning step are characterized by the ImageNet features. Therefore, we adopt this philosophy to initialize our visual grammar model ($\phi$ with $\Theta_{0}$) by pre-training the visual grammar model on the ImageNet dataset. Then, during CSS training, we will progressively train our visual grammar model at each learning step as aforementioned.
@@ 4.1 Implementations and Evaluation Protocols
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-101 100-150 all Major Minor
@@
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-101 100-150 all Major Minor
@@
Implementation Following common practices [19, 52, 3], we adopt DeepLab-V3 [6] with ResNet-101 [24] and SegFormer [56] with MiT-B3 [56] in our experiments. For the Visual Grammar model, we adopt the design of [35] with $L_{\phi}=12$ blocks of multi-head self-attention layers. The feature vectors from the last layer of the decoder are used for our $\mathcal{L}^{\alpha}_{Cont}$ loss. The value $\alpha$ is set individually for each dataset, i.e., $\alpha=5\times 10^{-2}$ for ADE20K, $\alpha=10^{-2}$ for VOC for Cityscapes. The details of our hyper-parameters are provided in the supplementary.
Table 2: Effectiveness of Scaling Factor $\alpha$ in $\mathcal{L}^{\alpha}_{Cont}$.
@@
Evaluation Protocols: We evaluate models on three standard datasets of CSS, i.e., ADE20K [61], Pascal VOC [22], and Cityscapes [15]. Following common practices [52, 4, 5], our experiments are conducted on the overlapped CSS settings. In particular, on ADE20K, we use three different settings, i.e., ADE20K 100-50 (2 steps), ADE20K 100-10 (6 steps), and ADE20K 100-5 (11 steps). On Pascal VOC, we evaluate FALCON in three benchmarks, i.e., VOC 15-5 (2 steps), VOC 15-1 (6 steps), and VOC 10-1 (11 steps). On Cityscapes, we conduct domain incremental experiments with three settings, i.e., Cityscapes 11-5 (3 steps), Cityscapes 11-1 (11 steps), and Cityscapes 1-1 (21 steps). Following [19, 4], the mean Intersection over Union (mIoU) metric is adopted in our comparison, including mIoU of the last learning step on initial classes, incremental classes, and all classes. In addition, to illustrate the fairness improvement, we report the mIoU of major and minor classes.
4.2 Ablation Study
Effectiveness of Fairness Contrastive Clustering Table 1 presents our results using DeepLab-V3 [7] with Resnet101 on ADE20K 100-50 and ADE20K 100-10 benchmarks. We evaluate the impact of the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}$ by comparing it with the vanilla contrastive clustering loss $\mathcal{L}_{Cont}$. As shown in our results, the overall performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. In addition, the fairness of the model has been promoted since the mIoU performance of major and minor groups was enhanced. We also study the impact of network backbones and cluster margin $\nabla$ in our supplementary.
Table 3: Effectiveness of Contributions of Our Proposed Losses.
@@ 4.2 Ablation Study
Effectiveness of Scaling Factor of Cluster Table 2 illustrates the experimental results of the impact of different scaling factor $\alpha$ on ADE20K 100-50 and Pascal VOC 15-5 benchmarks. As shown in Table 2, when the value of scaling factor $\alpha$ gradually decreases, the performance of our proposed approach is improved accordingly since the fairness contrastive loss in Eqn (4) tends to be more uniform across major and minor classes. However, when the scaling factor is too small ($\alpha=0.005$), the impact of the loss enforcement becomes weaker leading to the weaker enforcement of the fairness contrastive clustering, resulting in lower overall performance. In addition, we have observed that the higher the number of classes demands the higher the value of $\alpha$ since it will increase the compactness of more clusters.

### v3

@@ 3.3 An Efficient Unknown Class Modeling
An ideal CSS approach must be able to model the unknown classes without supervision, especially in open-set contexts [31, 63] where there could be multiple unknown classes or objects. Prior studies have adopted the pseudo-label strategies [21, 4] based on the model predictions to assign labels for seen classes while unseen classes have been ignored, thus resulting in non-discriminative features. [31, 63] improved the background modeling by using an additional prototypical representation for unknown classes. However, these approaches consider different unknown classes as one (i.e., $N_{U}=1$) resulting in non-distinguished representations of different unknown classes. Thus, modeling function $\phi$ in Eqn. (2) without supervision of different unknown classes (i.e., $N_{U}>1$) is challenging.
Although modeling $\phi$ to determine the single feature $\mathbf{f}$ belonging to the cluster $\mathbf{c}$ is challenging, prior studies in clustering [42, 73, 72] have suggested that determine a set of features $\{\mathbf{f}^{t}_{i}\}_{i=1}^{M}$ belonging to cluster $\mathbf{c}$ should be easier.
Figure 5: The Proposed Visual Grammar Model.
@@
This derives from the fact that even though the feature representations of different classes are different, the distributions of features around its cluster (termed as Visual Grammar) in the feature space should be similar among classes or clusters. As a result, by learning the distribution of features and their clusters, the model $\phi$ can determine whether a feature belongs to a cluster. Then, by learning the model $\phi$ on prior known clusters and features, the knowledge of $\phi$ can be adaptively applied to unknown clusters. Fig. 5 illustrates our visual grammar model of the cluster distributions.
@@
Limitations of Prior Clustering Methods. The traditional methods in clustering, e.g., KNN or density-based clustering [23], remain limited to noisy features leading to producing the incorrect cluster assignment. Meanwhile, the modern clustering methods, e.g., Graph Neural Networks (GNNs) [73, 72], require a large memory to build the affinity graph for clusters. In addition, GNNs often learn the local structures of graphs (or clusters) and accumulate them via the aggregation layers. Hence, the global structures of the clusters, i.e., visual grammar, are not well modeled by GNNs [42]. Therefore, to address these limitations, we introduced a new Attention-based Visual Grammar approach to efficiently model the distribution of features and their clusters via the self-attention mechanism [66].
@@
Remark 1: Given a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}$ where $\mathbf{f}^{\mathbf{c}}_{i}$ denotes the feature $\mathbf{f}_{i}$ belonging to the cluster $\mathbf{c}$, and $\forall i\in[1..M-1]:\cos(\mathbf{f}_{i}^{\mathbf{c}},\mathbf{c})\geq\cos(\mathbf{f}_{i+1}^{\mathbf{c}},\mathbf{c})$ the Visual Grammar of the cluster $\mathbf{c}$ parameterized by $\Theta$ can be defined as in Eqn. (5).
@@
$\footnotesize\begin{split}\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{f}^{\mathbf{c}}_{1},\mathbf{f}^{\mathbf{c}}_{2},...,\mathbf{f}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\\ \Leftrightarrow\Theta^{*}&=\arg\min_{\Theta}\mathbb{E}_{\mathbf{c},\{\mathbf{f}^{\mathbf{c}}_{i}\}_{i=1}^{M}}\left[-\log p(\mathbf{\Delta}^{\mathbf{c}}_{1},\mathbf{\Delta}^{\mathbf{c}}_{2},...,\mathbf{\Delta}^{\mathbf{c}}_{M},\mathbf{c},\Theta)\right]\end{split}$ (5)
@@
where $\Delta^{\mathbf{c}}_{i}=\mathbf{f}^{\mathbf{c}}_{i}-\mathbf{c}$. Eqn. (5) defines the visual grammar of the cluster by modeling the feature distribution of $\mathbf{f}_{i}^{\mathbf{c}}$ and its cluster center $\mathbf{c}$. Let $\phi:\mathbb{R}^{(M+1)\times D}\to[0,1]^{M}$ be a function receiving a center $\mathbf{c}$ and a set of $M$ features $\{\mathbf{f}_{i}\}_{i=1}^{M}$ $\left(\cos(\mathbf{f}_{i},\mathbf{c})\geq\cos(\mathbf{f}_{i+1},\mathbf{c})\right)$ to determine whether $\mathbf{f}_{i}$ belonging to $\mathbf{c}$, i.e., $\mathbf{u}=\phi(\mathbf{\Delta}_{1},\mathbf{\Delta}_{2},...,\mathbf{\Delta}_{M},\mathbf{c})$ where $\mathbf{\Delta}_{i}=\mathbf{f}_{i}-\mathbf{c}$, $\mathbf{u}=[u_{1},u_{2},...,u_{M}]$ and $u_{i}=1$ denotes $\mathbf{f}_{i}$ belong to cluster $\mathbf{c}$ and vice versa. Hence, the visual grammar model in Eqn. (5) can be modeled by the network $\phi$ as follows with parameter $\Theta$ as in Eqn. (6).
@@
Eqn. (6) aims to model the distribution of features around its cluster by learning the correlation of relatively topological structures $\mathbf{\Delta}_{i}$ of features $\mathbf{f}_{i}$ around cluster $\mathbf{c}$. Then, based on knowledge of the cluster distribution, the model $\phi$ is able to determine whether a feature $\mathbf{f}_{i}$ belongs to cluster $\mathbf{c}$. Hence, it is essential that the model $\phi$ has the ability to exploit the correlation between features $\mathbf{f}_{i}$ and cluster $\mathbf{c}$ to learn the topological structure of visual grammar. Therefore, we adopt the self-attention mechanism [66, 42] to efficiently model these feature correlations. Particularly, the model $\phi$ is formed by $L_{\phi}$ blocks of self-attention as follows,
$\footnotesize\begin{split}\mathbf{z}_{0}&=\operatorname{LN([\mathbf{\Delta}_{1},...,\mathbf{\Delta}_{M},\mathbf{c}])}+\boldsymbol{\beta},\quad\mathbf{a}_{l}=\mathbf{z}_{l}+\operatorname{MHSA}(\mathbf{z})\\ \mathbf{z}_{l+1}&=\mathbf{a}_{l}+\operatorname{MLP}(\operatorname{LN}(\mathbf{a}_{l})),\quad\quad\quad\quad\mathbf{u}=\operatorname{Proj}(\mathbf{z}_{L_{\phi}})\end{split}$ (7)
@@
Cluster Assignment via Visual Grammar. Instead of assigning the clusters based on the model prediction [21, 4, 46] or nearest cluster [63, 31] that are less effective, the cluster assignment in our approach will be performed by the visual grammar model, i.e., the visual grammar model will consider the $M$ closest features around cluster $\mathbf{c}$ to assign the cluster for these features. Then, the cluster assignments are used to compute our Fairness Contrastive Clustering loss. In addition, following common practices [21, 4, 46], we improve background shift modeling by using the cluster assignments of features as the pseudo labels of pixels.
Unknown Cluster Initialization. Prior work [63, 31] initialized a single unknown cluster ($N_{U}=1$), thus resulting in producing non-discriminative class-wise features. However, there should be more than a single unknown cluster ($N_{U}>1$) to produce discriminative features for different unknown classes. Therefore, our approach first initializes a list of potential unknown clusters at each learning step via DB-SCAN [23] on the features of unknown classes extracted by the current CSS model. For the new known class $\mathcal{C}^{t}$, we initialize these clusters based on the mean of their feature representations. Meanwhile, the clusters of known classes learned in previous steps are maintained. Further details can be found in the appendix.
@@ 3.4 Continual Learning Procedure
Fig. 3 illustrates the training procedure of our continual learning approach. At each learning step $t$, the CSS model $F$ with $\theta_{t}$ is trained with the Fairness Contrastive Clustering loss defined in Eqn. (4) and the previous visual grammar model $\phi$ with $\Theta_{t-1}$. In addition, we introduce a cluster regularizer $\mathcal{R}_{C}$ to avoid the clusters of different classes collapsing into a single cluster. Therefore, the entire CSS learning objective in our approach can be formed as in Eqn. (8).
$\scriptsize\vspace{-3mm}\begin{split}\arg\min_{\theta_{t}}\mathbb{E}_{\mathbf{x}^{t},\mathbf{\hat{y}}^{t}}\Big{[}\mathcal{L}_{CE}\left(\mathbf{y}^{t},\mathbf{\hat{y}}^{t}\right)+\lambda_{CL}\sum_{\mathbf{c}_{i}}\mathcal{L}^{\alpha}_{Cont}\left(\mathbf{F}^{t},\mathbf{c}_{i}\right)+\lambda_{C}\mathcal{R}_{C}(\mathbf{c})\Big{]}\end{split}$ (8)
@@
Training Procedure of Visual Grammar Model. At CSS learning step $t$, we adopt the visual grammar model trained on the previous learning step, i.e., $\phi$ with $\Theta_{t-1}$, to perform the cluster assignment for the contrastive clustering loss defined in Eqn. (2). Then, the visual grammar model $\phi$ with $\Theta_{t}$ at learning step $t$ will be learned (initialized from $\Theta_{t-1}$) on the features extracted from the dataset and the set of known clusters $\mathbf{c}$ up to the current learning step. Following [42], we sample a center $\mathbf{c}$ from the known clusters and its $M$ closest features to train the visual grammar model.
Initial Visual Grammar Model. At the first learning step $t=1$, since no clusters have been learned initially, the visual grammar model $\phi$ with $\Theta_{0}$ is unavailable. However, as common practices in CSS [21, 4, 5], the segmentation model is typically trained from a pre-trained backbone on ImageNet [19]. As a result, the features extracted at the first learning step are characterized by the ImageNet features. Therefore, we adopt this philosophy to initialize our visual grammar model ($\phi$ with $\Theta_{0}$) by pre-training the visual grammar model on the ImageNet dataset. Then, during CSS training, we will progressively train our visual grammar model at each learning step as aforementioned.
@@ 4.1 Implementations and Evaluation Protocols
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-100 100-150 all Major Minor
@@
$\mathcal{L}_{Cont}$ $\mathcal{L}^{\alpha}_{Cont}$ 0-100 100-150 all Major Minor
@@
Table 2: Effectiveness of Visual Grammar.
@@
0-100 101-150 all Major Minor
Nearest Cluster 44.3 11.5 33.4 51.5 24.3
Fixed $\phi$ 44.6 17.6 35.6 52.0 27.4
Adaptive $\phi$ 44.6 24.5 37.9 52.1 30.8
(b) ADE20K 100-10
0-100 101-150 all Major Minor
Nearest Cluster 40.1 14.3 31.5 48.7 22.9
Fixed $\phi$ 43.0 18.5 34.9 50.6 27.0
Adaptive $\phi$ 44.4 20.4 36.4 51.8 28.7
Implementation. Following common practices [21, 63, 3], we adopt DeepLab-V3 [7] with ResNet-101 [27] and SegFormer [70] with MiT-B3 [70] in our experiments. For the Visual Grammar model, we adopt the design of [42] with $L_{\phi}=12$ blocks of multi-head self-attention layers. The feature vectors from the last layer of the decoder are used for our $\mathcal{L}^{\alpha}_{Cont}$ loss. The value $\alpha$ is set individually for each dataset, i.e., $\alpha=5\times 10^{-2}$ for ADE20K, $\alpha=10^{-2}$ for VOC for Cityscapes. The details of our hyper-parameters are provided in the appendix.
Table 3: Effectiveness of Scaling Factor $\alpha$.
(a) ADE20K 100-50
$\alpha$ 0-15 16-20 all Major Minor
@@
Table 4: Effectiveness of Our Proposed Losses.
@@
Evaluation Protocols. We evaluate models on three standard datasets of CSS, i.e., ADE20K [77], Pascal VOC [24], and Cityscapes [17]. Following common practices [63, 4, 5], our experiments are conducted on the overlapped CSS settings. In particular, on ADE20K, we use three different settings, i.e., ADE20K 100-50 (2 steps), ADE20K 100-10 (6 steps), and ADE20K 100-5 (11 steps). On Pascal VOC, we evaluate FALCON in three benchmarks, i.e., VOC 15-5 (2 steps), VOC 15-1 (6 steps), and VOC 10-1 (11 steps). On Cityscapes, we conduct domain incremental experiments with three settings, i.e., Cityscapes 11-5 (3 steps), Cityscapes 11-1 (11 steps), and Cityscapes 1-1 (21 steps). Following [21, 4], the mean Intersection over Union (mIoU) metric is adopted in our comparison, including mIoU of the last learning step on initial classes, incremental classes, and all classes. In addition, to illustrate the fairness improvement, we report the mIoU of major and minor classes.
4.2 Ablation Study
Effectiveness of Fairness Contrastive Clustering. Table 1 presents our results using DeepLab-V3 [8] with Resnet101 on ADE20K 100-50 and ADE20K 100-10 benchmarks. We evaluate the impact of the fairness contrastive clustering loss $\mathcal{L}^{\alpha}_{Cont}$ by comparing it with the vanilla contrastive clustering loss $\mathcal{L}_{Cont}$. As shown in our results, the overall performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. In addition, the fairness has been promoted due to the mIoU improvement of minor groups. We also study the impact of network backbones and cluster margin $\nabla$ in our appendix.
Effectiveness of Scaling Factor of Cluster. Table 3 illustrates the experimental results of the impact of different scaling factor $\alpha$ on ADE20K 100-50 and Pascal VOC 15-5 benchmarks. As shown in Table 3, when the value of scaling factor $\alpha$ gradually decreases, the performance of our proposed approach is improved accordingly since the fairness contrastive loss in Eqn (4) tends to be more uniform across major and minor classes. However, when the scaling factor is too small ($\alpha=0.005$), the impact of the loss enforcement becomes weaker leading to the weaker enforcement of the fairness contrastive clustering, resulting in lower overall performance. In addition, we have observed that the higher the number of classes demands the higher the value of $\alpha$ since it will increase the compactness of more clusters.
Effectiveness of Loss Contributions. Table 4 illustrates the contributions of proposed learning objectives. For the model without using visual grammar, we only use a single unknown cluster ($N_{U}=1$) and adopt the nearest cluster strategies to assign clusters of unknown pixels. By using only cross-entropy loss, the mIoU performance remains low due to catastrophic forgetting and background shift problems. Meanwhile, with our fairness clustering loss $\mathcal{L}^{\alpha}_{Cont}$, visual grammar model $\phi$, and the cluster regularizer $\mathcal{R}$, the mIoU performance has been significantly improved to $37.9\%$ and $36.4\%$ on ADE20K 100-50 and ADE20K 100-10, respectively. Moreover, our FALCON has significantly promoted the fairness of segmentation models illustrated by the mIoU improvement of major and minor groups.
Effectiveness of Visual Grammar. We evaluate FALCON under three settings, i.e., Nearest Cluster, Fixed $\phi$ pretrained ImageNet (without updating on each learning step), and Adaptive $\phi$ (with updating on each learning step). As in Table 2, the mIoU result using only the nearest cluster remains ineffective. Meanwhile, the adaptive visual grammar model updated at each learning step further boosts the mIoU performance and promotes fairness, i.e., increased by $4.5\%$ and $4.9\%$ on ADE20K 100-50 and ADE20K 100-10 compared to the nearest cluster approach. In addition, we study the impact of choosing the number of features $M$ in the visual grammar model in our appendix. Fig. 6 illustrates the feature distributions of unknown classes (future class). As a result, our FALCON approach is able to model features of unknown classes into different clusters and produce better and more compact clusters compared to the one without Fairness Learning via Contrastive Attention.